A quality-control engineer of a high-tech company noted that 4% of all components provided by the company’s main supplier were of inferior quality (nonconforming items). Out of a sample of three hundred components provided by the supplier, what is the probability that more than 6% are defective?

0.0901

0.0764

0.0384

0.0708

None of the above
Answer #1

Solution

Given that,

p = 0.04

1 - p = 1-0.04=0.96

n = 300

= p =0.04

=  [p( 1 - p ) / n] = [(0.04*0.96) /300 ] = 0.01131

P( > 0.06) = 1 - P( <0.06 )

= 1 - P(( - ) / < (0.06 -0.04) /0.01131 )

= 1 - P(z < 1.77)

Using z table

= 1 -0.9616

=0.0384

probability=0.0384